OpenStudy (anonymous):

0.4p+8.1=4.3p-7.2-0.9p solve?

OpenStudy (anonymous):

0.4p + 8.1 = (4.3-0.9)p -7.2 0.4p + 8.1 = 3.2p - 7.2 15.3 = 3.6p p = 15.3/3.6 = 4.25